Westerville Company reported the following results from last year’s operations:

Sales $ 1,000,000
Variable expenses 300,000
Contribution margin 700,000
Fixed expenses 500,000
Net operating income $ 200,000
Average operating assets $ 625,000


At the beginning of this year, the company has a $120,000 investment opportunity with the following cost and revenue characteristics:

Sales $ 200,000
Contribution margin ratio 60 % of sales
Fixed expenses $ 90,000

The company’s minimum required rate of return is 15%.

8. If the company pursues the investment opportunity and otherwise performs the same as last year, what turnover will it earn this year? (Round your answer to 2 decimal places.)

10-a. If Westerville’s chief executive officer will earn a bonus only if her ROI from this year exceeds her ROI from last year, would she pursue the investment opportunity?

  • Yes

  • No

10-b. Would the owners of the company want her to pursue the investment opportunity?

  • Yes

  • No

    14. If Westerville’s chief executive officer will earn a bonus only if her residual income from this year exceeds her residual income from last year, would she pursue the investment opportunity
  • Yes

  • No

  • 15-a. Assume that the contribution margin ratio of the investment opportunity was 50% instead of 60%. If Westerville’s chief executive officer will earn a bonus only if her residual income from this year exceeds her residual income from last year, would she pursue the investment opportunity?
  • Yes

  • No

  • 15-b. Would the owners of the company want her to pursue the investment opportunity?

  • Yes

  • No
Answer #1

Total Sales = 1,000,000+200,000 = $1,200,000

Turnover Ratio= Sales/average operating assets

= (1,000,000+200,000)/(625000+120000)

= 1.6107 times

10.ROI last year = Net Operating Income/Average Operating assets

= 200,000/625000

= 32%

After investment = (200,000+30000)/745000

= 30.87%

No

10-b Yes, since the ROI is higher than the minimum required return

14.Yes

Since residual income from new investment = 30,000 – 120,000*15%

= $12,000

15-a.

Residual income from new investment = 200,000*50% -90,000 – 120,000*15%

= -$8000

NO

15-b NO